what percent of the total surface area of the oil container fill 4 inches up and the height is 10 inches with a radius of 6 inches?
Answer by addingup(3677)

Volume = r^2Pi*h
36*3.1416*10= 1130.97
36*3.1416*4= 452.39
452.39/1130.97= 0.40 or 40%